Consider the circuit below where: ΔΚǐat,,- 20V, Δν¡at,2= 90V, Ri-9000, R2 = 500Ω, and R3 = 130Ω R1 R3

How much voltage is dropped across the second resistor?
Answer #1

Writing KVL loop equation

let current in R1 be I1 ,in R2 be I2 and in R3 be I3

I1=I2+I3

20-90=900I1+500I2=>-70=1400I2+900I3.................(1)

90=130I3-500I2...........................(2)

solving 1 and 2

I2=-0.1425A

I3=0.1439 A

I1=I2+I3=0.00149 A

Voltage drop acrossR2=I2*R2=71.25


answered by: ANURANJAN SARSAM